# Richard A Berk

### v2

@@ 9 A Replication Study in Phoenix, Arizona
There are at least three reasons to undertake a replication in a new setting. First, the forecasting results just reported capitalize on extensive data exploration that can lead to cherry picked results; post-model-selection inference can be concern regardless of our efforts made to respond to the problem. Second, our findings acquire far more credibility and usefulness if they apply to areas in addition to the Pacific Northwest. That credibility and usefulness can affect subject-matter considerations and pressing policy concerns. Finally, other geographical areas bring not just different climate scenarios, but different data and different data problems. Especially important is to determine if our results are robust to new data flaws and research design constraints.
@@
To address these matters of generalizability, we obtained AIRS daily data for the Phoenix area for 63 grid cells from 2019 through 2023. Data were not available for 2024. The measured variables were the very same used in the PNW analysis. However, the topography in the greater Phoenix area is not just very different from the PNW, but far more homogeneous. One might wonder, for instance, if a quasiresonant explanation applies.
We began with a simple and ambitious aspiration. Using a crossover design much like before, training with random forests on the same variables used previously, and with similar algorithmic methods to that used for the PNW forecasts, could we reproduce the PNW results in Phoenix? An exact replication would be effectively impossible to produce, but could we get results properly to inform affirmative judgments about reproducibility?
@@
There were immediate and obvious design issues. We had about half the number of grid cells for the Phoenix area than for the Pacific Northwest. There was also a substantial number of grid cells with missing data. Between the two, we anticipated a much smaller number of observations overall. One result could much weaker power and even more sparsity. We needed a way to increase the number of usable grid cells within in the intent of our earlier crossover design.
@@
The Phoenix heat dome arrived at the beginning of July 2023, not at end of June, 2021, as it had in the Pacific Northwest. For an alternative crossover design, we proceeded the much as before, but the faux heat waves A and B for ordinary summer temperatures were for July of 2021 and July of 2022 respectively, before the actual Phoenix extreme heat wave. This means that the predictors lagged 14 days earlier would fall toward the middle of June for 2021, 2022 and 2023. For all three Phoenix years, we used a binary response variable coded “1” for the heat dome in July of 2023, and “0” for no heat dome in July of 2021 and July 2022.222222 We could have included a greater number of years back in time, but all would for faux July heat waves. The number of appearances made by each grid cell could be increased along with the overall number of observations. But we were concerned that the binary response variable would become badly unbalanced.
@@
The coding is consistent with that done for the PNW forecasting analysis. All three time periods were then stacked to make single data set; each grid cell in principle was observed in three different years and in principle there would be 189 observations representing one extreme heat wave and two faux heat heat waves. The crossover design would again provide sufficient control for temporally stable features of grid cells.
Figure 7: The True Heat Wave is compared simultaneously both to Faux Heat Waves A and B. Each has a binary post-test, and predictors 14 days earlier. The True Heat Wave began the first day July, 2023. The two faux heat waves began on the first day of July 2021 and the first day of July 2022. Were we trying to be fully consistent with notation used in the research design for the PNW, the time indices start at t0 on 6/1/2021, and the time index values for the true heat wave and the 2022 faux heat wave would be much larger.
@@
But the potential problem from missing data remained. When we removed all rows (i.e., grid cells) with any missing data, we lost 186 grid cells from the total 189. Clearly this was disastrous. Fortunately, when the included variables were limited to just those selected for the PNW forecasts, the number of grid cells fell to “only” 128. Because the data were lost through cloud cover and occasional downtime on the remote sensing satellite, one can argue that the data were missing in blocks conditionally at random, and exchangeability still holds. Nevertheless, the large amount of missing data raises concerns about the representativeness of the Phoenix data and about some instabilities that could be introduced. The design we used is shown in Figure 7.
@@
We were pleasantly surprised. The error rate for Phoenix was a little under 4%, roughly comparable to the error rate for the PNW. As before, all three predictors made important contributions to the fit although their rank order differed. The atmospheric water vapor mass mixing ratio was now the top ranked. A different importance ordering was anticipated because the topography and climate in the Phoenix area is so very different from the topography and climate in the Pacific Northwest.
@@
Despite great differences in climate and topography, Figure 6 for the PNW and Figure 8 for the greater Phoenix area have their smoothers showing much the same S-Shaped relationship for atmospheric temperature at an altitude of 8. Both relationships begin flat for the lowest values before the slope turns positive, and end flat at the highest values. The Phoenix relationship is shifted a bit to the right in response to its higher temperatures. It also somewhat less smooth because of data that are more sparse.
@@
Comparing the PNW and Phoenix smooths for tropopause height, both start out flat before turning positive. For the PNW, the overall pattern is roughly linear and positive thereafter. For the greater phoenix area, the relationship turns negative at around 15,500 meters.
@@
For the water vapor mixing ratio, neither the smoother for the PNW nor for the greater Phoenix area show easily interpreted relationships. This helps explain why in both geographical settings, the water vapor mix contributes little to the fit. The good news is that it seems not to be needed. In both settings the fit is excellent, and the two variables largely responsible have somewhat similar relationships with the presence of an extreme heat wave. Finally, we again employed nested conformal prediction sets. As before, forecasts were very precise (i.e., one outcome class chosen) even with coverage probabilities of .90.
It is important to emphasize that while the PNW analysis was steeped in extensive data exploration, the Phoenix analysis had none. There was no cherry picking of results whatsoever. Yet classification accuracy was the nearly the same for both and two of the three partial dependence plots had much the same look and feel despite some differences in detail.
10 Crossover Design Balance and Problems from Endogenous Sampling
@@
Recall that the crossover designs for the PNW and for the greater Phoenix area began by constructing the data for a true and a faux heat waves with the same grid cells in each; each had the same number of observations. This led to excellent balance that remedied the rare events problem. However, it is challenging to think of a naturally occurring population of grid cells in which such balance could have occurred. For example, if the population were comprised of all grid cells in the pacific northwest for each day during June of 2021, for about 13% of the days, a grid cell would be subject to an extreme heat wave. Yet, in our designs, for approximately 33% or 50% of the days, an actual heat wave would have produced extreme surface temperatures. The marginal distribution of the curated 0/1 response variable was no longer distributed like the 0/1 response variable for the month of June. Our curation process oversampled extreme heat wave days, which is a form of endogenous sampling (because the response variable is treated as endogenous).
@@
As noted earlier, there is a literature on problems that can result when data collection alters the prior distribution of the response variable compared to its distribution in the population of interest (Manski and Lerman, 1977; Manski and McFadden, 1981). For example, one might over-sample cancer cases that would otherwise be too small a fraction of the data for sufficient pattern recognition accuracy (Wang et al., 2021). But the price is empirical results that can lack many desirable statistical properties. There can be, for instance, badly distorted estimates of fitted probabilities.
For discrete response variables, several remedies have been proposed (Cosslett, 1993; Waldman, 2000). An entire dataset can be weighted in one of several ways so that the endogenous (response) variable once again can provide a consistent estimate of the response variable’s distribution in the target population. Sometimes the weighting is readily apparent. Sometimes it is buried in the statistical machinery along with performance enhancements.
@@
Appropriate weighting is easily undertaken.232323 We take a transparent approach because we have seen no formal proofs of the more elaborate proposed remedies for random forests, gradients boosting, or nonparametric classifiers more generally, when used for forecasting. For example, whether or not there is an extreme heat wave is the endogenous (response) variable in Figure 7. Suppose $P(heat)$ is the probability of an extreme heat wave in the target population. $P^{\ast}(heat)$ is the probability of an extreme heat wave in the curated data on hand. The two probabilities are different because of the manner in which the data collected and organized. There is a pair of remedial weights (Manski and Lerman, 1977; Manski and McFadden, 1981). For each grid cell $i$ that experienced an extreme heat wave, $weight_{heat,i}=\frac{P(heat_{i}=1)}{P^{\ast}(heat_{i}=1)}$. For each grid cell $i$ that did not experience an extreme heat wave, $weight_{noheat,i}=\frac{P(heat_{i}=0)}{P^{\ast}(heat_{i}=0)}$. These weights can be incorporated easily within a weighting option commonly available in machine learning software.242424 As before, a heat wave is coded equal to “1.” The absence of a heat wave is coded equal to “0.”
@@
There are several difficulties with such solutions. First, in many applications it is not apparent what the proper population is or even should be. An earlier example was whether the population for studying a rare extreme heat wave should be daily ARIS data for the most recent 5 years, daily AIRS data for the most recent year, or daily AIRS data for the most recent month. The choice can depend heavily how heat wave forecasting will be undertaken in practice. If essential responses to extreme heat waves are determined month to month, monthly data might be preferred. But then, a trained forecasting algorithm for each month would be needed, perhaps updated year to year. The work load could be reduced if forecasters were prepared to bet that rare extreme head waves only occur in the summer months. The forecasters would then be backing into the kinds of designs we used in the PNW and greater Phoenix area.
@@
Second, for rare discrete events, the proper weighting proposals often recommended resurrect the very problem that the oversampling of rare events tried to remedy; it is not clear that the proposed remedies anticipated forecasting very rare events. The oversampled rare events are made empirically rare once again and preferred machine learning software may not run at all, or may run but give consequential warming messages. For example, the fitted probabilities may fall outside of the 0 to 1 range.
@@
Third, Therneau and Atkinson (2023) have made clear that for tree-based classifiers such as classification trees, the prior distribution of the response is related to the ratio of the number of false negatives to the number of false positives (or the reciprocal). That ratio is often used to represent the relative costs of false negatives to false positive (i.e., all costs, not just monetized costs). For example, if for a binary response variable, a classifier produces 40 false positives and 20 false negatives, there are 2 false positives for every false negative. This can be interpreted to mean that false negatives are twice as costly as false positives, which may help explain why the classifier fits the data so that there are fewer of them; one false negative is “worth” two false positives.
More fundamentally, because the false negatives and false positives can have different costs, the algorithm is minimizing risk, not loss, a formulation convincingly introduced in the classic book by Luce and Raiffa (1957). It is no longer clear how endogenous sampling problems apply because the prior distribution of the response variable and differential losses for false negatives and false positives are related, and together they affect the risk. The underlying statistical machinery might be more appropriately built on asymmetric risk functions applicable to algorithmic forecasting (Ridnick et al., 2021). In fact, the entire discussion of unbalanced classification problems might be usefully reframed to make the consequences of false negatives and false positives a matter for asymmetric loss functions to address.252525 The ratio of false negatives to false positives is something about which stakeholders might care deeply. How much more (or less) costly a failure to correctly forecast an extreme heat wave (i.e., a false negative) may be compared to an extreme heat wave false alarm (i.e., a false positive). This is not a statistical decision. It is a policy decision in which subjective values and judgments are essential. Ideally, a preferred target ratio of the number of false negatives to the number of false positives is determined before the analysis begins and is then used to help formulate the data collection and key features of the forecasting algorithm. This has worked well in other applied settings (Berk, 2012).
@@
The conflation of corrections for sample imbalance with the estimated misclassification cost ratio is easily illustrated. Using the expressions above for the weights, the PNW forecasting exercise was repeated. Recall that the response variable was binary, and there were the same number of grid cells that experienced a heat wave as did not experience a heat wave. Using the month of June, 2021 as the target population, for a little more than 13% of the days, the 144 grid cells experienced the extreme heat wave. For the other days, the 144 grid cells did not. In the data analyzed, for 50% of the days, the 144 grid cells experienced the extreme heat wave. A difference between 13% and 50% is the kind of disjuncture addressed by Manski and Lerman (1977). Note the substantial disparity materialized even though by choosing June, 2021 and using grid cells as the study units, the imbalance was statistically manageable; the extreme heat wave was not disastrously rare.
@@
Before weighting, the classification error rate was a little over 5%, and there were nearly the same number of false negatives as false positives. The cost ratio was about 1.0, conveying equal costs. When Manski-Lerman weighing was applied, the classification error rate rose to nearly 14%, and the cost ratio of false positives to false negatives increased from 1.0 to 9.67. False positives were now well over 9 times more costly than false negatives. In other words, the costs of heat wave false alarms were substantially up-weighted compared to the costs of true alarms not given. Many stakeholders might strongly prefer the reverse. The statistical point is that changing the prior of the response variable can dramatically change the relative costs of false negatives and positives provided by a classifier. To concentrate only on corrections for unrepresentative, response variable priors is far too narrow.
We currently favor using, when feasible, data collection and curation methods that alter the prior joint distribution of the heat wave data, as we did for the PNW and greater Phoenix area. Random forests, gradient boosting and other classifiers can respond as if the costs of false negatives and false positives are asymmetric. However, the cost ratio should be determined by stakeholders as the research is being planned. We currently are not operating under those circumstances. Without weighting, the classification cost ratios for the forecasting reported above were about 1 to 1, which may be a useful place to start the discusion. Results for other cost ratios can be computed with weighting or by altering directly the proportion of heat wave observations in the data through selective case augmentations of, or deletions from, the study dataset. The more general approach is to employ asymmetric loss functions directly, a discussion of which would take as far afield.

### v3

@@ 9 A Replication Study in Phoenix, Arizona
Repeating a forecasting exercise in a new geographical setting can have several benefits. First, one is able to consider whether the statistical approaches employed earlier generalize well. Does the formulation of a particular data generating process more broadly apply? Are gain scores instructive elsewhere? Are the random forests and the GA genetic algorithm appropriate in other locations? Second, there is substantive generalizability. In particular, using the same variables with their values realized elsewhere, how well reproduced is the confusion table from the Pacific Northwest? All of these questions can be properly addressed with no feature engineering, no “bake-offs” between algorithms, no model selection, no p-hacking, and no filtering of reported results.
9.1 Data Challenges
@@
We obtained AIRS daily data for the Phoenix area for 63 grid cells from 2019 through 2023. Data were not available for 2024. The measured variables were the very same used in the PNW analysis. However, the climate and topography in the greater Phoenix area is very different from the PNW and far more homogeneous. The Pacific Northwest is lush. The greater Phoenix area is largely a desert. We picked the greater Phoenix area precisely because of the strong replication challenges it posed.
@@
We had about half the number of grid cells for the Phoenix area than for the Pacific Northwest. There were also a substantial number of grid cells with missing data. Between the two, we anticipated a much smaller number of observations overall. One result could much weaker statistical power coupled with still more sparsity. We needed a procedure to increase the number of usable grid cells for the within-subject (i.e., grid cells) approach.
@@
According to many media accounts, the Phoenix heat dome arrived at the beginning of July 2023, not at end of June, 2021, as the extreme heat wave had in the Pacific Northwest. We proceeded much as before, but faux heat waves A and B, representing ordinary summer temperatures, were positioned in early July of 2021 and early July of 2022 respectively, well before the actual Phoenix extreme heat wave. The predictors lagged 14 days earlier fell toward the middle of June in 2021, 2022 and 2023. For all three Phoenix years, we used a binary response variable coded “1” for the extreme heat wave in early July of 2023, and “0” for faux extreme heat waves in early July of 2021 and early July 2022. The actual heat wave apparently began the around first day July, 2023. Correspondingly, the two faux heat waves began on the first day of July 2021 and the first day of July 2022.
@@
All three time periods were then stacked to make single data set. Each grid cell in principle was observed in three different years and in principle, there would be 189 observations overall, each exposed to one set of extreme heat wave precursors and two sets of faux heat wave precursors. One benefit is that the collection of study subjects was the same under all three conditions; the could be no changes in sample composition.
Figure 7: Research Design Configurations for the Greater Phoenix Forecasts
@@
But the potential problems from missing data remained. When we included only the three precursor variables required for comparability that were used for the PNW forecasts, the number of grid cells with complete data fell to 128. Because the data were lost through cloud cover and occasional downtime on the remote sensing satellite, one might argue that the data were missing in blocks conditionally at random. Nevertheless, the large amount of missing data raises concerns about the representativeness of the Phoenix data and about substantial instabilities that could be introduced. There would be further complications should there be differential amounts of non-random missingness when the grid cells were exposed to true heat wave precursors versus the when the grid cells were exposed to the faux heat wave precursors. The design structure is shown in Figure 7.171717 Because the data were likely generated in an unknown, spatially dependent matter, all of the missing data imputation methods with which were familiar seem inappropriate.
Figure 8: fig: Phoenix Histograms for Predictors, Actual Heat Wave in red, faux heat waves ind blue, and their overlap in purple.
9.2 Comparisons Between True Heat Wave and Faux Heat wave Predictor Histograms
As before, we compare the distributions for each of three predictors when exposed to the true heat wave precursors compared to the faux heat wave precursors. The amount of missing data is far more pronounced for the true heat wave data. We expected about 33% of the data to be from the faux heat wave cases, and about 67% of the cases to be from the faux heat wave cases. After listwise deletion, 27% of the observations were from the true heat wave, and 73% were from the faux heat waves. The numbers of observations were 35 and 93 respectively. True heat wave cases are noticeably underrepresented, and the sample size is small enough to be worrisome.
For the true heat wave and faux heat waves, Figure 8 shows the top two plots different predictor distributions for tropopause height and for temperature at level 8. But compared to Phoenix, the true heat wave predictor values tend to be toward the left of the faux heat wave values. The bottom plot for the water vapor mixing ratio, is again inconclusive. Compared to the PNW data, one might speculate that the larger fraction of missing values from the true heat wave precursors are disproportionately from greater tropopause heights and higher temperatures. That might account for some results that at this point seem counterintuitive.
9.3 Classification Results
We began simply. Using the random forests classifier already trained for with data from the Pacific Northwest, we obtained the fitted probabilities with input data from the greater Phoenix area. That is, the trained algorithm was fixed. The data from Phoenix were employed solely to construct fitted probabilities. Whatever the relationships captured for the PNW were being implicitly assumed to apply in Phoenix.
The resulting Phoenix confusion table has an overall misclassification proportion of a little over .16, which is about 3 times larger the overall misclassification proportion for the PNW. There are no false positives, but the false negative fraction is 60%. This substantial difference in the proportion of false negatives compared to the proportion of false positives may be in part a result of the marginal distribution of the response variable. Whereas there are for the PNW data nearly the same number of actual heat wave precursor exposures as faux heat wave precursor exposures, the proportions for Phoenix have approximately two faux heat wave exposures for each actual heat wave exposure. With weighting methods described in the next section, the empirical ratio of false negatives to false positives could be made to approximate 1 to 1, but the complications that can follow would need to be addressed.
A second comparability assessment was undertaken using the random forest classifier just as before with the Phoenix data structured as in Figure 7. All three three predictors used in the PNW analysis are used again. After deletion of all rows with any missing data, there were 128 observations. With so few observations in each, substantial instability was expected. The implicit assumption was that the same precursors would matter, but probably not in the same way as before.
We were pleasantly surprised by the confusion table produced. The overall misclassification rate for Phoenix is a little over .03, about the same as error rate for the PNW. There were no false positives, and a false negative rate of approximately .11 The difference was likely explained in part by the relative decline the number of in the heat wave cases, now accounting for just a little over a quarter of the data. All three predictors contribute substantially to the fit, although the role of atmospheric temperature 14 days earlier is more dominant than for the PNW.181818 Given the likely instability because of the modest number of training data observations, we examined the confusion tables following from 10 additional random splits of the data. Overall misclassification proportions ranged from a little over .05 to a little under .12. Often there were no false positives. By most standards, classification accuracy was quite good but lacked the stability desired.
@@ 9.3 Classification Results
Despite the histogram differences just described and great differences in climate and topography, Figure 6 for the PNW and Figure 9 for the greater Phoenix area have their smoothers showing much the same S-shaped relationship between the fitted probabilities and atmospheric temperature at an altitude of 8.191919 As before, the vertical variation in the plots is visually attenuated. Both relationships begin flat for their lowest values before their slopes turns positive, and end flat at the highest values. The Phoenix relationship is shifted a bit to the right in response to its higher temperatures. It also somewhat less smooth because of data that are more sparse.
@@
Comparing the PNW and Phoenix smooths for tropopause height, both start out flat before turning positive. For the PNW, the overall pattern is roughly linear and positive thereafter. For the greater phoenix area, the relationship smoothly turns negative at around 15,500 meters, and the tropopause heights start as somewhat higher values.
@@
For the water vapor mixing ratio at altitude 8, neither the smoother for the PNW nor for the greater Phoenix area show easily interpreted relationships. This helps explain why in both geographical settings, the water vapor mix by itself contributes relatively little to the fit. The good news is that it seems not to be needed.
The apparent contradiction between the predictor comparative histogram outcomes and the partial dependence plots is perhaps explained by the difference between marginal and conditional distributions. The partial dependence plots are conditioned so that adjustments are made for the dependence between the three predictors. For example, in the Phoenix data, the correlation between temperature at level 8 and tropopause height is .83.
Despite the excellent classification accuracy, one might wonder if there remain dependencies in the Phoenix analysis. Given the spatial nature of the data, one might be curious about spatial dependence between grid cells implying relationships not yet addressed. Classifier residuals were computed. The spatial correlogram R procedure gstat, authored and maintained by Edzer Pebesma, was applied. Figure 10 shows the result. The correlogram shows that the measure of association, here Moran’s I, fluctuates around 0.0 and is not systematically related to the distance between spatial residuals.202020 The residuals can in principle range between ±$\pm 1.0$. The discrete unit for the plot is .01. The horizontal axis is in distance multiples of .01. After a Bonferroni correction for multiple testing, for none of the associations is the null hypothesis of no association rejected. The nominal critical values was .05.
Figure 10: Spatial Correlogram at Different Distances in Units of .01 with Statistical Tests for Each Subject to the Bonferroni Correction for the Nominal Critical Value of .05.
It is important to emphasize that although the PNW analysis was steeped in extensive data exploration, the Phoenix analysis had none. Yet classification accuracy is similar for both, and two of the three partial dependence plots are roughly similar in both locations. Moreover, one might think that for Phoenix, conformal prediction sets can now be used. As before, however, we knew the days on which the Phoenix heat wave occurred before the data analysis began and as before, grid cells in advance were coded homogeneously. When grid cells were exposed to true heat wave precursors, they were all code equal to 1.0. When grid cells were exposed to faux heat wave precursors, they were all code equal to 0.0. The data are neither IID nor exchangeable.
10 Problems from Endogenous Sampling
@@
The cross-over designs for the PNW and for the greater Phoenix area led to excellent balance that remedied the rare events problem for our binary outcome of heat wave or not. However, it is challenging to think of a naturally occurring population of grid cells that could be a plausible source of the data in which such balance would have occurred. Our curation process substantially oversampled extreme heat wave days, which is a form of troublesome endogenous oversampling.
@@
As noted earlier, there is a literature on problems that can result when data collection alters the prior distribution of the response variable compared to its distribution in the population of interest (Manski and Lerman, 1977; Manski and McFadden, 1981). For example, in a study of cancer detection in lung X-rays, one might oversample cancer cases compared to cancer-free cases. Absent the oversampling, there would be too few “positives” for a sufficiently accurate detection assessment (Wang et al., 2021). But the price is badly distorted estimates of the sensitivity and specificity for the cancer diagnoses.
For numeric response variables, clever sampling strategies have been proposed that, in effect, oversample for extreme values (Ragone et al., 2018). Endogenous sampling is explicit and worrisome. For discrete response variables, our focus here, several related remedies have been proposed (Cosslett, 1993; Waldman, 2000). A dataset can be weighted in different ways so that the imbalance is corrected and the endogenous (response) variable once again can provide a valid estimates. Sometimes the weighting is buried in the statistical machinery along with other performance enhancements, but it is still altering the sample imbalance. For example, when Matthew’s correlation, also know as the phi coefficient, is used as an objective function for a classifier to maximize, it can help correct imbalance (Dourard it al., 2024). But because the cost ratio target of false positives to false negative is set to 1.0, the marginal distribution is, in effect, being re-weighted.
@@
A simple form of appropriate weighting is easily undertaken. For example, whether or not there is an extreme heat wave can be the endogenous (response) variable as shown in Figure 7. Suppose $P(heat)$ is the probability of an extreme heat wave in the target population, and $P^{\ast}(heat)$ is the probability of an extreme heat wave in the curated data on hand. The two probabilities are different because of the manner in which the data were collected and organized. There is a pair of remedial weights (Manski and Lerman, 1977; Manski and McFadden, 1981). For each grid cell $i$ that experienced an extreme heat wave, $weight_{heat,i}=\frac{P(heat_{i}=1)}{P^{\ast}(heat_{i}=1)}$. For each grid cell $i$ that did not experience an extreme heat wave, $weight_{noheat,i}=\frac{P(heat_{i}=0)}{P^{\ast}(heat_{i}=0)}$. These weights can be incorporated easily within a weighting option commonly available in machine learning software.212121 As before, a heat wave is coded equal to “1.” The absence of a heat wave is coded equal to “0.”
@@
There are several difficulties with such solutions. To begin, in many applications it is not apparent what the proper population is or even should be. An earlier example was whether the population for studying a rare extreme heat wave should be daily ARIS data for the most recent 5 years, daily AIRS data for the most recent year, or daily AIRS data for the most recent month. The choice can depend heavily how heat wave forecasting will be undertaken in practice. If essential responses to extreme heat waves are determined month to month, monthly data might be preferred. But then, a trained forecasting algorithm for each month would be needed, perhaps updated year to year. The work load could be reduced if forecasters were prepared to assume that rare extreme head waves only occur in the summer months. The forecasters would then be backing into the kinds of designs we used in the PNW and greater Phoenix area.
@@
In addition, for rare discrete events, the weighting proposals typically recommended can restore the very problem that the initial oversampling of rare events tried to remedy. Their compensatory down-weighting makes the formerly rare events rare once again. Preferred machine learning software may not run at all, or may run with unsettling warning messages. For example, the fitted probabilities may fall outside of the 0 to 1 range.
@@
Some adjustments back toward the target population can trigger deeper difficulties. Therneau and Atkinson (2023) show that for tree-based classifiers such as classification trees, the prior distribution of the response is related to the ratio of the number of false negatives to the number of false positives (or the reciprocal). That ratio is often used to represent the relative costs of false negatives to false positive (i.e., all costs, not just monetized costs). For example, if for a binary response variable a classifier produces 40 false positives and 20 false negatives, there are 2 false positives for every false negative. This can be interpreted to mean that false negatives are twice as costly as false positives, which may help explain why the classifier fits the data so that there are relatively fewer of them; one false negative is “worth” two false positives. The costs are asymmetric.
In response to asymmetric costs, there are weighting adjustments leading to what Kuhn and Johnson (2013: Chapter 6) call “cost sensitive training.” Related thinking can be found in focal loss approaches (Lin et al., 2017) and implicitly in gradient boosting (Friedman, 2001). Rednick and his colleagues (2021) offer the most involved approach that weights more heavily the hard to classify observations, and a specified outcome class to can be given a higher priority as well.222222 We coded their loss function in keras3 in R. It ran, but did far worse that the usual cross-entropy loss, perhaps because the classifier was already too successful with the within-subject design and the balanced data that resulted. There were too few false negatives and false positives usefully to weight. We also found tuning a conceptual challenge when we tried to return to a highly imbalanced source population response variable and with a plausible stakeholder consensus cost ratio. In their classic book, Luce and Raiffa (1957) formulate the problem as risk minimization rather than loss minimization. Broadly framed, asymmetric loss functions are required to account for asymmetric fitting costs.
@@
The conflation of corrections for sample imbalance with the estimated misclassification cost ratio is easily illustrated. Using the expressions above for the weights, the PNW forecasting exercise was repeated. Recall that the response variable was binary, and there were the same number of grid cells that experienced a heat wave as did not experience a heat wave. Using the month of June, 2021 as the target population, for a little more than 13% of the days, the 144 grid cells experienced the extreme heat wave. For the other days, the 144 grid cells did not. In the data analyzed, for 50% of the days, the 144 grid cells experienced the extreme heat wave. A disparity between 13% and 50% may be more extreme than anticipated by Manski and Lerman (1977), but we can proceed for illustrative purposes.
@@
After weighting, the population imbalance was approximately reproduced. But that’s not all. Before weighting, the classification error rate was a little over 5%, and there were nearly the same number of false negatives as false positives. The cost ratio was about 1.0, conveying equal costs. When Manski-Lerman weighing was applied, the classification error rate rose to nearly 14%, and the cost ratio of false positives to false negatives increased from about 1.0 to nearly 10.0. False positives were now almost 10 times more costly than false negatives. In other words, the costs of heat wave false alarms were substantially up-weighted compared to the costs of true alarms not given. Many stakeholders might strongly prefer the reverse. In short, weighting changes the prior distribution of the response variable and the relative costs of false negatives to false positives.
For the forecasting results reported for the PNW and the greater Phoenix area, we had no access to stakeholders nor any concrete policy context. The numbers of false negatives and false positives were very few and of roughly equal size. Classifier training determined an empirical cost ratio of about 1 to 1. Going forward in a real policy setting combined with stakeholder input, the within-subject research design could be revised to better capture practical concerns. The main requirement is that grid cells continue to capitalize on the within-subject structure. Additional faux heat waves could easily be added so that a more appropriate, but manageable, imbalance was achieved. Another possibility might be to train the classifier on balanced data but to improve fitted values, use as input an altered dataset with many more faux heat waves.